OpenStudy (anonymous):

log(x-30)=??? please help with procedure

OpenStudy (mathmale):

State your goal, please. You're given log(x-30)=??? Note that the base of this log is 10. Given log(x-30)=??? how would you eliminate the "log" operator? Best to write this as y=log(x-30) and then perform the same operations on both sides.

OpenStudy (anonymous):

sorry the equation was log(x-30)=3-log^x log base 10

OpenStudy (anonymous):

please help :( thank you so much